# Estonian WinoGrande Dataset: Comparative Analysis of LLM Performance on Human and Machine Translation

### v1

@@ 6.3.  The Impact of Machine Translation
To study the effect of machine translation on the results, we had a qualified translator (the same person who did the human translation) manually analyze the subset of 1,617 semantically comparable samples (i.e., those that did not undergo localization, adaptation or correction) to identify instances where the machine translation had altered the meaning of the original sentence. In this labeling, meaning was the sole evaluation criterion; grammatical issues, unnatural phrasing, or other surface-level errors were not considered problematic unless they affected the sentence’s meaning. The guiding question was: Does the machine-translated schema convey the same meaning as the original English schema?. [^5] These annotations will be released with the dataset.
@@
We then compared results on the subset of schemas where machine translation had altered the meaning to those where machine translation had retained the meaning (MT columns in Table 6). The shifts in meaning has a clear impact on performance—with one exception (Qwen 2.5-72B), all models perform worse on machine translated schemas with altered meaning. With the simple prompt, the difference in accuracies between the averages of the two subsets is 4.7% and with the detailed prompt it is 7.2%. However, it must be noted that, besides human-perceived shift in meaning, machine translated instances contain other artifacts that may affect the accuracy even when the meaning is retained. Thus, we also compared subset of schemas with retained meaning with the human translated version of the same subset (rows in the lower part of Table 6). This comparison shows that the accuracies on the machine translated datasets are lower (3.8% with the simple prompt and 3.3% with the detailed prompt) indicating that there are issues beyond altered meaning in the machine translated data that also potentially affect task integrity.

### v2

@@ 6.3.  The Impact of Machine Translation
To study the effect of machine translation on the results, we had a qualified translator (the same person who did the human translation) manually analyze the subset of 1,617 semantically comparable samples to identify instances where the machine translation had altered the meaning of the original sentence. In this labeling, meaning was the sole evaluation criterion; grammatical issues, unnatural phrasing, or other surface-level errors were not considered problematic unless they affected the sentence’s meaning. The guiding question was: Does the machine-translated schema convey the same meaning as the original English schema?. [^4] These annotations will be released with the dataset.
@@
We then compared results on the subset of schemas where machine translation had altered the meaning to those where machine translation had retained the meaning (MT columns in Table 8). The shifts in meaning have a clear impact on performance—with one exception (Qwen 2.5-72B), all models perform worse on machine translated schemas with altered meaning. With the simple prompt, the difference in accuracies between the averages of the two subsets is 4.7% and with the detailed prompt it is 7.2%. However, it must be noted that, besides human-perceived shift in meaning, machine translated instances contain other artifacts that may affect the accuracy even when the meaning is retained. Thus, we also compared the subset of schemas with retained meaning with the human translated version of the same subset (rows in the lower part of Table 8). This comparison shows that the accuracies on the machine translated datasets are lower (3.8% with the simple prompt and 3.3% with the detailed prompt) indicating that there are issues beyond altered meaning in the machine translated data that also potentially affect task integrity.
@@ 7.  Discussion
We noted a substantial difference in the number of items the annotators labeled as undecidable. Although the discrepancy is numerically large, we do not consider it significant, as the annotators’ accuracy remained close to the reported human performance level (94%) on the English test set Sakaguchi et al. (2019).
@@ Acknowledgements
Appendix A Translation Prompts
A.1.  Simple Translation
{promptbox}
A.2.  Detailed Prompt for Single Items
{promptbox}
A.3.  Translation Prompt for Twin Sentences
{promptbox}
The following are sentence pairs with a gap expressed with the underscore and two answer options. Your task is to translate each sentence pair into Estonian so that it reads as if originally written in Estonian: natural, fluent, and culturally appropriate.
The sentence pairs have a lexical overlap of at least 70%, and their translations must also maintain at least 70% overlap. This overlap should be preserved consistently, without introducing unnecessary variation. Do not vary style, synonyms, or sentence structure between them. Answer options must also remain exactly the same in these sentence pairs.
Do not translate word for word. Always choose the contextually correct meaning of each word, especially adjectives, idioms, and figurative expressions — never use a literal dictionary equivalent if it does not fit the situation. Ensure that both answer options remain grammatically and semantically valid in the sentences, without revealing or implying which is correct. Pay particular attention to number and case: both answer options must agree in grammatical form so they can fit the same syntactic slot.
If the two answer options translations differ in grammatical number (singular vs plural), replace one option with a contextually appropriate alternative so that both options agree in form (e.g. both plural nouns, both singular nouns). The replacement must still preserve the fairness and neutrality of the question.
Do not fill the gap. Replace all names with randomly chosen ones from the provided table, matching gender and applying correct Estonian case endings. Both names in a sentence must be the same gender, and names may be reused across sentence pairs but not within the same sentence. For sentence pairs, use the same names consistently in both sentences.
Do not add names where none exist. Maintain neutrality and difficulty: the question must stay fair, challenging, and unbiased. You will receive a JSON list containing two examples as input. Your output should also be JSON list containing the translated sentences and answer options. Make sure you translate only the values and not the keys. Please output only the JSON list, nothing else. Do not escape with backticks or add addtional line breaks.